# Racka: Efficient Hungarian LLM Adaptation on Academic Infrastructure

### v1

@@ Abstract
We present Racka, a lightweight, continually pretrained large language model designed to bridge the resource gap between Hungarian and high‑resource languages such as English and German. Racka employs parameter‑efficient continual pretraining via Low‑Rank Adaptation (LoRA) on a Qwen‑3 4B backbone, making the recipe practical on A100 (40GB)‑based HPC clusters with low inter‑node bandwidth. To better match the training distribution, we replace and adapt the tokenizer, achieving substantially improved tokenization fertility for Hungarian while maintaining competitive performance in English and German. The model is trained on 160B subword tokens drawn from a mixture of internet and high‑quality curated sources, with a composition of 44% Hungarian, 24% English, 21% German, and 11% code. This data mix is chosen to mitigate catastrophic forgetting and preserve high‑resource language capabilities during continual pretraining. Our preliminary results indicate modest but stable results in language adaptation.
@@ 1.1 The Challenge of Linguistic Digital Sovereignty for Medium-Resource Languages
The proliferation of Large Language Models (LLMs) has marked a transformative era in artificial intelligence, yet this progress has been unevenly distributed across the world’s languages (ali2024surveylargelanguagemodels). The current landscape is dominated by models pretrained on vast corpora composed predominantly of English and a few other high-resource languages, creating a significant performance and resource disparity for less-resourced linguistic communities (zhong2025opportunitieschallengeslargelanguage). For medium-resource languages such as Hungarian, a Finno-Ugric language characterized by its agglutinative nature and rich morphology, this gap is particularly pronounced. Off-the-shelf multilingual models often exhibit suboptimal performance due to insufficient representation in training data and tokenizers that are ill-suited to their linguistic structure. This is particularly the case for open-source models, which visibly struggle with Hungarian grammar.
This disparity underscores a critical challenge to linguistic digital sovereignty: the capacity for a linguistic community to develop, deploy, and benefit from AI technologies that are not only proficient in its language but are also attuned to its unique cultural and contextual nuances. Achieving this requires the creation of models trained on nationally relevant data and, crucially, developed using methodologies that are accessible to local research communities, who may not have access to the hyperscale computing infrastructure available to large industrial labs, making researchers rely on shared HPC infrastructure.
@@ 1.2 Core Contributions
In response to these challenges, we introduce Racka, a 4-billion-parameter language model for Hungarian adapted trough a pragmatic and resource-efficient approach from a Qwen3-4B reasoning model. The model’s name pays homage to the Hortobágy Racka, an indigenous Hungarian sheep breed and llama alternative. Racka was trained using publicly accessible infrastructure—Hungary’s Komondor supercomputer—rather than state-of-the-art proprietary hardware.
This paper presents a comprehensive recipe for language adaptation that is both effective and practical for research communities with access to national HPC clusters. Our primary contributions are threefold:
@@
  2. 2.
    Optimized Tokenization for Hungarian: We underscore the critical role of tokenizer adaptation for morphologically rich, agglutinative languages. We document the design, training, and evaluation of a new tokenizer that dramatically improves tokenization efficiency (i.e., subword fertility) for Hungarian, while carefully preserving the performance of the base model on high-resource control languages like English and German.
  3. 3.
    Curation of a Large-Scale Multilingual Corpus: We describe the assembly and curation of a 160 billion subword token dataset representing state-of-the-art size in training a Hungarian-centric language model. We justify its strategic composition (44% Hungarian, 43% high-resource languages, and 11% code) as a pragmatic approach to mitigate catastrophic forgetting during the continual pretraining process.
@@ 2.1 Paradigms of Language Model Adaptation
Adapting existing foundation models to new languages, domains, or tasks is a central challenge in modern NLP. The prohibitively high cost of training LLMs from scratch has spurred research into more efficient adaptation strategies. (csaki2023efficientlyadaptingpretrainedlanguage; khade2024challengesadaptingmultilingualllms) As an industry standard practice for language adaptation, continual pretraining (CPT) is widely adopted. This method focuses on further self-supervised training on large, unlabeled corpora to expand a model’s foundational knowledge. (vanroy2024geitje7bultraconversational; laion2024leolm; ke2023continual)
The work on Racka falls squarely within the CPT paradigm, specifically targeting what the literature defines as “Language Expansion” and “Domain Adaptation”. This framing positions our effort not as the creation of a final, task-specific model, but as the development of a more capable foundational model for Hungarian, which can subsequently be adapted for downstream tasks. A primary obstacle in any continual learning scenario is catastrophic forgetting, where a model’s performance on previously learned tasks or languages degrades significantly as it learns new information (gupta2023continual). A simple yet effective strategy to mitigate this is data replay, which involves mixing data from the original training distribution with the new data. Our multilingual data composition, which retains a significant portion of English and German text, is a direct application of this principle. Other alternatives include careful train scheduling, intense regularizations, or the analysis of task-vector shifts. We refrain from these methods in favor of data replay, as it directly controls our training objective and does not require extensive experimentation (li2024examiningforgettingcontinualpretraining). Previous Hungarian language adaptation work also applied this approach successfully (szentmihalyi2024pretraining; csaki2024sambalingoteachinglargelanguage; gyozo2025pulillumix).
@@ 2.2 Parameter-Efficient Methods for Continual Pretraining
The computational burden of updating the billions of parameters in an LLM has led to the development of Parameter-Efficient Fine-Tuning (PEFT) methods. These techniques aim to adapt models by training only a small fraction of their total parameters, drastically reducing memory and compute requirements. Among the most widely adopted PEFT methods is Low-Rank Adaptation (LoRA). (hu2021loralowrankadaptationlarge) LoRA operates on the hypothesis that the change in model weights during adaptation has a low intrinsic rank. This aligns well with our aim to train on limited-memory GPUs with low throughput internode connections. By keeping the base model frozen, it is shown to inherently aid in overcoming catastrophic forgetting of the base model’s knowledge thus making them ideal for CPT (li2024examiningforgettingcontinualpretraining).
@@ 2.3 The Landscape of Hungarian Language Models
  • NYTK Models: The Hungarian Research Centre for Linguistics (NYTK) has been a key contributor, developing a series of models under the “PULI” family. These include models in the range of 7-8B parameters, both monolingual and continually pretrained. Latest models include instruction-tuned and multilingual (English, Chinese) models. (yang-puli-gptrio; gyozo2025pulillumix; yang-llumix-llama)
  • OTP Bank Models: A major industrial and government supported initiative resulted in the OTP 1.5B and OTP 13B bilingual models, which were trained on an extensive dataset in collaboration with SambaNova Systems. These models demonstrated the feasibility of developing high-performance models for Hungarian at a large scale. (szentmihalyi2024pretraining)
  • International Initiatives: The ecosystem also includes models adapted from multilingual open-source foundations. SambaLingo-Hungarian-Base, for instance, adapted a Llama-2 7B model by training on 59 billion tokens from the Hungarian portion of the Cultura-X dataset. (csaki2024sambalingoteachinglargelanguage) Similarly, projects like OpenEuroLLM have fine-tuned models like Gemma for improved Hungarian performance. Other European initiatives such as EuroLLM and Salamandra include Hungarian in their supported languages. (openeurollm2025hungarian; eurollm2025; gonzalezagirre2025salamandratechnicalreport)
This existing landscape highlights a clear and valuable niche for the Racka project. While previous efforts have focused on larger models (OTP-13B), monolingual training (PULI-GPT-3SX), or adaptation on smaller datasets (SambaLingo), Racka’s contribution is its unique combination of a lightweight (4B) backbone, and a continual pretraining on an exceptionally large multilingual corpus (160B tokens), and a resource-efficient methodology explicitly designed for and validated on publicly accessible national HPC infrastructure. One key insight tha can be accrued from the papers above is that for smaller languages, continual pretraining of already capable models results in better performning LLMs than training new models from scratch.
@@ 3.1 Corpus Design and Composition
The foundation of any successful language model is the quality and scale of its training data. For Racka, we curated a comprehensive multilingual corpus totaling approximately 160 billion subword tokens with a dataset size of $639$ GB. The composition of this corpus was a strategic design choice aimed at achieving robust language adaptation for Hungarian while simultaneously mitigating the effects of catastrophic forgetting on the backbone model’s existing capabilities. The final data mixture, detailed in Table 1, allocates the majority of the data to our target language, Hungarian, while retaining substantial portions of high-quality English, German, and source code data. The inclusion of English and German serves as a form of explicit data replay to preserve the model’s proficiency in these high-resource languages, as these are the most commonly utilized foreign languages in Hungary. The code component was included to maintain the model’s strong reasoning and structured-text processing abilities, which are known to be beneficial for a wide range of downstream analytical tasks (ma2024at).
@@ 3.2 Data Sources
The corpus was assembled from a combination of large-scale web crawls and high-quality curated datasets to ensure both breadth and quality. The primary source for web data across all languages was the Common Crawl repository, with all data sourced from crawls conducted prior to our cutoff date of October 2024. (commoncrawl2025)
For English, we sample approximately 3838B tokens sourced from non-web text-based subsets of The Pile (gao2020pile800gbdatasetdiverse), and the heavily filtered fineweb corpus (penedo2024finewebdatasetsdecantingweb). We sample approximately 3434B tokens of German text from the similarly pre-filtered occiglot-fineweb (occiglotfineweb2024) and 1818B tokens of code from The Stackv2 (lozhkov2024starcoder2stackv2), slightly oversampling structured text formats and markups to better represent them compared to code files. 7070B Hungarian tokens were collected from various sources including Common Crawl, academic repositories, court rulings, movie subtitles, Wikipedia and digital news articles. Following Nemeskey:2020, the web and news subcorpora were heavily filtered and deduplicated on both the URL and document-level. Paragraphs that appeared frequently across the respective subcorpora were also discarded. The aggregated per-language corpus sizes are reported in Table 1; the composition of the Hungarian corpus is detailed in Appendix B. Note that the fertility of our adapted tokenizer is smaller for Hungarian, thus we produce less tokens in Hungarian compared to the other subsets.
@@ 4.1 Backbone Model Architecture
The foundation for our work is the Qwen-3 4B model, an open-source transformer-based LLM developed by Alibaba Cloud accessed through Hugging Face [^1] https://huggingface.co/Qwen/Qwen3-4B. We selected this model as our backbone due to its strong baseline performance, favorable open-source license (Apache 2.0), and an architecture well-suited for our objectives. The Qwen-3 4B model features $4$ billion total parameters ($3.6$ billion non-embedding) distributed across $36$ transformer layers (yang2025qwen3technicalreport).
Key architectural specifications include the use of Grouped Query Attention (GQA) for efficient inference, the SwiGLU activation function for improved performance, and RMSNorm for stable training. (yang2025qwen3technicalreport) The model supports a native context length of 32,768 tokens, making it capable of processing long documents without truncation, which could be extended via interpolation of the positional encoding. (su2023roformerenhancedtransformerrotary) We specifically chose the reasoning and instruction-tuned variant of the model, hoping that its inherent capabilities would provide a stronger starting point and ultimately benefit in downstream performance after our continual pretraining phase. To our knowledge, this makes Racka the first reasoning model to be adapted to Hungarian.
@@ 4.2 Tokenizer Adaptation for Hungarian
A key hypothesis of our work is that the default tokenizer of a multilingual model is suboptimal for a morphologically rich, agglutinative language like Hungarian. Such tokenizers, trained on a distribution dominated by English, tend to fragment Hungarian words into numerous, less meaningful subword units, leading to longer sequence lengths, increased computational cost, and potentially poorer model performance (csaki2023efficientlyadaptingpretrainedlanguage). Several previous models have also experimented with tokenizer adaptation. To address this issue, we developed a new, adapted tokenizer for Racka, slightly modifying previously used adaptation algorithms. The process involved several steps:
@@
    We then merged this new Hungarian-specific vocabulary with the original vocabulary of the Qwen-3 tokenizer. Non-latin character-based multi-byte tokens were removed, and novel tokens were inserted resulting in an adapted vocabulary that retained all of the original latin alphabet tokens while adding 32 00032\ 000 new, Hungarian-optimized tokens learned from the Hungarian-only tokenizer. See Appendix A for details.
@@
    The model’s architecture was updated to accommodate this change: specifically, the token embedding matrix and the language-model head (the final output layer) were altered to match the new tokenizer. To provide a strong initialization for the newly added tokens in the embedding matrix, we employed the VIPI (Vocabulary Initialization with Partial Inheritance) vocabulary-transfer technique (mosin2022finetuning). For each new token, we used the original tokenizer to break it down into its constituent subwords. The new token’s initial embedding vector was then calculated as the average of the embedding vectors of these constituent subwords from the original, pretrained embedding matrix. This approach provides a semantically meaningful starting point for the new tokens, facilitating faster convergence during training.

### v2

@@ Abstract
We present Racka-4B, a lightweight, continually pretrained large language model designed to bridge the resource gap between Hungarian and high‑resource languages such as English and German. Racka employs parameter‑efficient continual pretraining via Low‑Rank Adaptation (LoRA) on a Qwen3-4B backbone, making the recipe practical on A100 (40GB)‑based HPC clusters with low inter‑node bandwidth. To better match the training distribution, we replace and adapt the tokenizer. By swapping 32k tokens, we achieve 47% lower tokenization fertility and generation latency for Hungarian while maintaining competitive performance in English and German. The model is trained on 160B subword tokens (85B whitespace words) drawn from a mixture of internet and high‑quality curated sources, with a composition of 44% Hungarian, 24% English, 21% German, and 11% code to mitigate catastrophic forgetting. Our evaluations indicate modest but stable results in language adaptation with similar LM-Eval-Harness results but increased overall performance on HULU and OpenHuEval compared to the original Qwen and the latest Puli models. The results also showcase that Racka-4B is capable of Hungarian chat with English reasoning even in the absence of explicit Hungarian post-training on these tasks.
@@ 1.1 The Challenge of Linguistic Digital Sovereignty for Medium-Resource Languages
The proliferation of Large Language Models (LLMs) has marked a transformative era in artificial intelligence, yet this progress has been unevenly distributed across the world’s languages (Ali and Pyysalo, 2024). The current landscape is dominated by models pretrained on vast corpora composed predominantly of English and a few other high-resource languages, creating a significant performance and resource disparity for less-resourced linguistic communities (Zhong et al., 2025). For medium-resource languages such as Hungarian, a Finno-Ugric language characterized by its agglutinative nature and rich morphology, this gap is particularly pronounced. Off-the-shelf multilingual models often exhibit suboptimal performance due to insufficient representation in training data and tokenizers that are ill-suited to language-specific morphology. This is particularly the case for open-source models, which visibly struggle with Hungarian grammar.
This disparity highlights a critical challenge to linguistic digital sovereignty. Linguistic communities need the capacity to develop, deploy, and benefit from AI technologies that are not only proficient in their language but are also attuned to their unique cultural and contextual nuances. Achieving this requires the creation of models trained on nationally relevant data and, crucially, developed using methodologies that are accessible to local research communities. Likely, most research groups will not have access to the hyperscale computing infrastructure available to large industrial labs and will have to rely on shared, and often underpowered, HPC infrastructure.
@@ 1.2 Core Contributions
In response to these challenges, we introduce Racka, a 4-billion-parameter language model for Hungarian adapted trough a pragmatic and resource-efficient approach from a Qwen3-4B reasoning model. The model’s name pays homage to the Hortobágy Racka, an indigenous Hungarian sheep breed and llama alternative. Racka was trained using publicly accessible infrastructure, Hungary’s Komondor supercomputer, rather than state-of-the-art proprietary hardware. Our primary contributions are as follows:
@@
    Curation of a Large-Scale Multilingual Corpus: We describe the assembly and curation of a 160 billion subword token dataset representing state-of-the-art size in training a Hungarian-centric language model. We justify its strategic composition (44% Hungarian, 43% high-resource languages, and 11% code) as a pragmatic approach to mitigate catastrophic forgetting during the continual pretraining process. The inclusion of English and German is justified by the census report of the Hungarian Central Statistical Office [^1] https://nepszamlalas2022.ksh.hu/eredmenyek/vegleges-adatok/kiadvany/, as these are the most common foreign languages in Hungary.
  2. 2.
    Optimized Tokenization for Hungarian: We underscore the critical role of tokenizer adaptation for morphologically rich. We document the design, training, and evaluation of a new tokenizer that dramatically improves tokenization efficiency (i.e., subword fertility) for Hungarian, while carefully preserving the performance of the base model on high-resource control languages like English and German.
  3. 3.
@@
  4. 4.
    The first Hungarian reasoning LLM: We present Racka, which to our knowledge is the first Hungarian LLM with reasoning output. The model is openly available on Hugging Face as elte-nlp/Racka-4B [^2] https://huggingface.co/elte-nlp/Racka-4B.
@@ 2.1 Paradigms of Language Model Adaptation
Adapting existing foundation models to new languages, domains, or tasks is a central challenge in modern NLP. The prohibitively high cost of training LLMs from scratch has spurred research into more efficient adaptation strategies (Csaki et al., 2023; Khade et al., 2025). As an industry standard practice for language adaptation, continual pretraining (CPT) is widely adopted. This method focuses on further self-supervised training on large, unlabeled corpora to expand a model’s foundational knowledge (Vanroy, 2024; Schuhmann et al., 2024; Ke et al., 2023).
The work on Racka falls squarely within the CPT paradigm, specifically targeting what the literature defines as “Language Expansion” and “Domain Adaptation”. This framing positions our effort not as the creation of a final, task-specific model, but as general fine-tuning of a foundational model for Hungarian, which can subsequently be adapted for downstream tasks. A primary obstacle in any continual learning scenario is catastrophic forgetting, where a model’s performance on previously learned tasks or languages degrades significantly as it learns new information (Ke et al., 2023). A simple yet effective strategy to mitigate this is data replay, which involves mixing data from the original training distribution with the new data. Our multilingual data composition, which retains a significant portion of English and German text, is a direct application of this principle. Other alternatives include careful train scheduling, intense regularizations, or the analysis of task-vector shifts. We refrain from these methods in favor of data replay, as it directly controls our training objective and does not require extensive experimentation (Li and Lee, 2024). Previous Hungarian language adaptation work also applied this approach successfully (Szentmihályi et al., 2024; Csaki et al., 2024; Yang et al., 2025d).
@@ 2.2 Parameter-Efficient Methods for Continual Pretraining
The computational burden of updating the billions of parameters in an LLM has led to the development of Parameter-Efficient Fine-Tuning (PEFT) methods. These techniques aim to adapt models by training only a small fraction of their total parameters, drastically reducing memory and compute requirements. Among the most widely adopted PEFT methods is Low-Rank Adaptation (LoRA; Hu et al., 2021). LoRA operates on the hypothesis that the change in model weights during adaptation has a low intrinsic rank. This aligns well with our aim to train on limited-memory GPUs with low throughput internode connections. By keeping the base model frozen, it is shown to inherently aid in overcoming catastrophic forgetting of the base model’s knowledge thus making it ideal for CPT (Li and Lee, 2024).
@@ 2.3 The Landscape of Hungarian Language Models
  • NYTK Models: The Hungarian Research Centre for Linguistics (NYTK) has been a key contributor, developing a series of models within the “Puli” family. These include models in the range of 7-8B parameters, both monolingual and continually pretrained. Latest models include instruction-tuned and multilingual (English, Chinese) models (Yang et al., 2023, 2025d, 2025c).
  • OTP Bank Models: A major government-supported initiative, lead by OTP Bank, demonstrated the feasibility of developing high-performance models for Hungarian at a large scale (Szentmihályi et al., 2024). The resulting models were trained on an extensive dataset on custom SambaNova hardware. Unfortunately, these models are not available to the public.
  • International Initiatives: The ecosystem also includes models adapted from multilingual open-source foundations. SambaLingo-Hungarian-Base, for instance, adapted a Llama-2 7B model by training on 59 billion tokens from the Hungarian portion of the Cultura-X dataset (Csaki et al., 2024). Similarly, projects like OpenEuroLLM have fine-tuned models like Gemma for improved Hungarian performance. Other European initiatives such as EuroLLM and Salamandra include Hungarian in their supported languages. (OpenEuroLLM, 2025; Martins et al., 2025; Gonzalez-Agirre et al., 2025)
This existing landscape highlights a clear and valuable niche for the Racka project. While previous efforts have focused on larger models (OTP-13B), monolingual training (PULI-GPT-3SX), multilingual training (PULI-GPTrio) or adaptation on smaller datasets (SambaLingo), Racka’s contribution is its unique combination of a lightweight (4B) backbone, continual pretraining on an exceptionally large multilingual corpus (160B tokens), and a resource-efficient methodology explicitly designed for and validated on publicly accessible national HPC infrastructure. One key insight that can be accrued from the papers above is that for smaller languages, continual pretraining of already capable models results in better performing LLMs than training new models from scratch.
@@ 3.1 Corpus Design and Composition
The foundation of any successful language model is the quality and scale of its training data. For Racka, we curated a comprehensive multilingual corpus totaling approximately 160 billion subword tokens with a dataset size of $639$ GB. The composition of this corpus was a strategic design choice aimed at achieving robust language adaptation for Hungarian while simultaneously mitigating the effects of catastrophic forgetting on the backbone model’s existing capabilities. The final data mixture, detailed in Table 1, allocates the majority of the data to our target language, Hungarian, while retaining substantial portions of high-quality English, German, and source code data. The code component was included to maintain the model’s strong reasoning and structured-text processing abilities, which are known to be beneficial for a wide range of downstream analytical tasks (Ma et al., 2024).
@@ 3.2 Data Sources
The corpus was assembled from a combination of large-scale web crawls and high-quality curated datasets to ensure both breadth and quality. The primary source for web data across all languages was the Common Crawl repository, with all data sourced from crawls conducted prior to our cutoff date of October 2024 (Common Crawl, 2025).
For English, we sample approximately 3838B tokens sourced from non-web text-based subsets of The Pile (Gao et al., 2020), and the heavily filtered FineWeb corpus (Penedo et al., 2024). We sample approximately 3434B tokens of German text from the similarly pre-filtered occiglot-fineweb (Occiglot Project, 2024) and 1818B tokens of code from The Stackv2 (Lozhkov et al., 2024), slightly oversampling structured text formats and markups to better represent them compared to code files. 7070B Hungarian tokens were collected from various sources including Common Crawl, academic repositories, court rulings, movie subtitles, Wikipedia and digital news articles. Following Nemeskey (2020), the web and news subcorpora were heavily filtered and deduplicated on both the URL and document-level. Paragraphs that appeared frequently across the respective subcorpora were also discarded. The aggregated per-language corpus sizes are reported in Table 1; the composition of the Hungarian corpus is detailed in Appendix B. Note that the fertility of our adapted tokenizer is smaller for Hungarian, thus we produce less tokens in Hungarian compared to the other subsets.
@@ 4.1 Backbone Model Architecture
The foundation for our work is the Qwen3-4B model, an open-source transformer-based LLM developed by Alibaba Cloud accessed through Hugging Face [^3] https://huggingface.co/Qwen/Qwen3-4B. We selected this model as our backbone due to its strong baseline performance, favorable open-source license (Apache 2.0), and an architecture well-suited for our objectives. The Qwen3-4B model features $4$ billion total parameters ($3.6$ billion non-embedding) distributed across $36$ transformer layers (Yang et al., 2025a).
Key architectural specifications include the use of Grouped Query Attention (GQA) for efficient inference, the SwiGLU activation function for improved performance, and RMSNorm for stable training (Yang et al., 2025a). The model supports a native context length of 32,768 tokens, making it capable of processing long documents without truncation, which could be extended via interpolation of the positional encoding (Su et al., 2024). We specifically chose the reasoning and instruction-tuned variant of the model, hoping that its inherent capabilities would provide a stronger starting point and ultimately benefit in downstream performance after our continual pretraining phase. To our knowledge, this makes Racka the first reasoning model to be adapted to Hungarian.
@@ 4.2 Tokenizer Adaptation for Hungarian
A key hypothesis of our work is that the default tokenizer of a multilingual model is suboptimal for a morphologically rich, agglutinative language like Hungarian. Such tokenizers, trained on a distribution dominated by English, tend to fragment Hungarian words into numerous, less meaningful subword units, leading to longer sequence lengths, increased computational cost, and potentially poorer model performance (Csaki et al., 2023). Several previous models have also experimented with tokenizer adaptation. To address this issue, we developed a new, adapted tokenizer for Racka, slightly modifying previously used adaptation algorithms. The process involved several steps:
@@
    We then merged this new Hungarian-specific vocabulary with the original vocabulary of the Qwen3 tokenizer. Rare, typically non-latin character-based multi-byte tokens were removed, and novel tokens were inserted resulting in an adapted vocabulary that retained all of the original latin alphabet tokens while adding 32 00032\ 000 new, Hungarian-optimized tokens learned from the Hungarian-only tokenizer. See Appendix A for details.
@@
    The model’s architecture was updated to accommodate this change: the token embedding matrix and the language-model head (the final output layer) were altered to match the new tokenizer. To provide a strong initialization for the newly added tokens in the embedding matrix, we employed the VIPI (Vocabulary Initialization with Partial Inheritance) vocabulary-transfer technique (Mosin et al., 2023). For each new token, we used the original tokenizer to break it down into its constituent subwords. The new token’s initial embedding vector was then calculated as the average of the embedding vectors of these constituent subwords from the original, pretrained embedding matrix. This approach provides a semantically meaningful starting point for the new tokens, facilitating faster convergence during training.